OpenStudy (anonymous):

Does the function y=8x represent a direct or inverse variation.

jimthompson5910 (jim_thompson5910):

this is direct variation because it is of the form y = k*x where k is some constant number

jimthompson5910 (jim_thompson5910):

other examples of direct variation y = 3x (in this case, k = 3) y = 19x (in this case, k = 19) y = (1/2)x (in this case, k = 1/2)

OpenStudy (anonymous):

Thank you.

jimthompson5910 (jim_thompson5910):

you're welcome